OpenStudy (anonymous):

Choose the graph which matches the following system of equations. 8x + 5y = -5 7x - 3y = 3

OpenStudy (anonymous):

We need the graphs.

OpenStudy (anonymous):

A. B. C. D.

OpenStudy (anonymous):

crap ! it wont post

OpenStudy (anonymous):

The best way to do this is just put both equations in slope-intercept form, graph them, and see which answer choice looks like the one you made.

OpenStudy (anonymous):

i dont know how ?

OpenStudy (anonymous):

hello.

OpenStudy (anonymous):

To put them in slope-intercept, rearrange them so they equal y. 8x + 5y = -5, 7x - 3y = 3. For the first one, subtract 8x over to get 5y=-5-8x, then divide by 5 to get y=8/5x-1. 8/5x is the slope. It means that from any point on the line, you can go up 8 units and right 5 to get another point on the line. Also, you can go down 8 and left 5. The -1 is the y-intercept. It means that -1 is where the line crosses the y axis. So you know that one of the points is (-1,0). Plot that point first, then go up 8 and right 5. Plot that point, then draw a line between them. Do exactly the same thing for the other equation. For that one, the slope-intercept form will be (7x-3y=3, -3y=3-7x, y=7/3x-1.) The S-I form is y=7/3x-1. For that one, the slope is 7/3x, so the next point on the line can be found by going up 7 and right 3. (IMPORTANT: this is a different line. Do not start and your last point from the other line.) For this one, the y-intercept is also -1. Since both of them have that, you know they will cross at (-1,0). Graph it the same way you did the other. Now, look at your answer choices. Pick the one that looks like what you just drew.

OpenStudy (anonymous):

I love you !!!